You have opened your own word-processing service. You bought a personal computer, and paid $5,000 for it. However, due to the cost changes in the computer industry, the current price of an equivalent machine is $2,500. You could sell any used machine for $1,000. If you were not word processing, you could earn $20,000 per year at an alternative job. Assume that the interest rate is 10%. You can also hire an assistant who can do everything that you can do for $20,000 per year (you would still continue to do word processing).

One person using one computer can produce 11,000 typed pages per year, and the price per page for your service is $2.

You are considering three options: (1) expand your business by hiring an assistant; (2) leave your business the way it is; (3) shut down. Based on the costs and revenues above, which should you do? Explain and show any relevant calculations.
1
Expert's answer
2017-05-25T11:33:10-0400
If you consider to expand your business by hiring an assistant, then your additional profit for the first year will be TP = TR - TC = 2*11,000 - (20,000 + 2,500) = -$500, but the next year you will earn additional profit of $2,000, so you should hire an assistant.